OpenStudy (anonymous):

The sum of two numbers is 72, and twice their difference is 24. What is the smaller of the two numbers?

myininaya (myininaya):

x+y=72 2(x-y)=24

OpenStudy (anonymous):

one number = x another - y now, just read the problem... x+y=72 2(x-y)=24 solve the system for z & y - your numbers

OpenStudy (anonymous):

I got 12 for x. Am I on the right track?

myininaya (myininaya):

x+y=72 x-y=12 2x=72+12=84 so x=84/2=42

myininaya (myininaya):

y=72-x x-(72-x)=12 x-72+x=12 2x=12+72 2x=84 x=42

OpenStudy (anonymous):

So, how would I find the other number? Plug it in?